OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x – 3y = −12 and passes through the point (−3, 4).

OpenStudy (anonymous):

first you have to change 4x-3y=-12 to y= mx+b. so: 4x-3y=-12 (-3y=-4x-12)/-3 y= 4/3 + 4 m= 4/3 then, using the point slope formula y-y1= m(x-x1): (-3,4) y-4= 4/3 [x-(-3)] y-4= 4/3 (x+3) 3(y-4)= 4(x+3) 3y-12 = 4x+12 -4x+3y= 24 is your answer